OpenStudy (anonymous):

Find f of g and g of f..find the domain of each function and each composition function.. f(x)=square root of x+5 g(x)=x^2

OpenStudy (anonymous):

fog(x)=f(g(x) \[=\sqrt{x ^{2}}+5=x+5\] gof(x)=g(f(x)) \[=(\sqrt{x}+5)^{2}\]

OpenStudy (anonymous):

can you show me the solution..so that i can answer # 2 and 3 questions.. pls :)

OpenStudy (anonymous):

First off, does f=\[\sqrt{x+5}\] or \[\sqrt{x}+5\]?

OpenStudy (anonymous):

\[\sqrt{x+5}\]

OpenStudy (anonymous):

fog(x) and gof(x) are the composite functions of f and g

OpenStudy (anonymous):

okay; f(g(x))= \[\sqrt{x^2+5}\] g(f(x))= x+5 the domain of f is x > or equal to -5 so that no negative under the root g's domain is all the real numbers f(g(x)) domain is all reals since x^2 always greater than zero same with g(f(x)) domain is all reals